A company has two departments, Y and Z that incur wage expenses. An analysis of the total wage expense of $34,000 indicates that Dept. Y had a direct wage expense of $5,000 and Dept. Z had a direct wage expense of $8,000. The remaining expenses are indirect and analysis indicates they should be allocated evenly between the two departments. Departmental wage expenses for Dept. Y and Dept. Z, respectively, are:

$15,500; $18,500.

$18,500; $15,500.

$17,000; $17,000.

$5,000; $8,000.

$10,500; $10,500.
Answer #1

Indirect wages = 34,000 - (5000+8000) = 21,000

Departmental wages for dept Y

= 5000 + (21,000/2)

= 15,500

Departmental wages for dept Z

= 8000 + (21,000/2)

= 18,500

Option A